# Projection Theorems in the Presence of Expansions

### v1

@@ 1. Introduction
Let us fix some notation in order to state the main results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
@@
The following is the main theorem in this paper.
@@
Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{3}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
@@
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}\cdot(\#F)$
@@
We also state a similar result in all dimensions $n\geq 3$, which relies on the aforementioned work of [GGW22].
Let $12\leq k\leq n-$ 1. For all $t\geq 0$ and all $r\in[0,1]$, let $\pi_{t,r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be
$\pi_{t,r}^{(k)}(w)=\Bigl{(}e^{t}w\cdot\xi^{\prime}(r),e^{t}w\cdot\xi^{\prime\prime}(r),\ldots,e^{t}w\cdot\xi^{(k-1)}(r),w\cdot\xi^{(k)}(r)\Bigr{)}$
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$ and
$\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}.$
Using [GGW22, Thm. 2.1] and some elementary arguments, one can establish the following
1.2 Proposition.
Let $0<\alpha\leq k$, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{n}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{n}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in F$ and all $\delta\geq\delta_{0}$}$
where $C_{0}\geq 1$.
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-\alpha^{\prime}t}\delta^{\alpha-\varepsilon}\cdot(\#F)$
where $\alpha^{\prime}=\alpha$ if $10<\alpha\leq k-$ 1 and $\alpha^{\prime}=\alpha-1$ otherwise.
It is worth noting that Proposition 1.2 does not imply Theorem 1.1; indeed $\alpha^{\prime}=\alpha-1<1/10$ for $101\leq\alpha<11/$ 10. That said, it seems plausible that the approach developed in [GGW22] can be modified to prove a stronger version of Proposition 1.2 which then would imply (and perhaps improve) Theorem 1.1. However, our proof of Theorem 1.1 relies on much softer arguments and is an adaptation of [OV18].
Also as previously mentioned, our precise formulations here (which at least at first glance may appear somewhat ad hoc) are motivated by recent applications in homogeneous dynamics [LM23, LMW22, LMWY23].
@@
Let $\mu$ denote the normalized counting measure on $F$:
@@
3. Proof of Proposition 1.2
Let us now turn to the proof of Proposition 1.2. As it was already mentioned, we will prove this theorem using [GGW22, Thm. 2.1]. Let us begin by recalling the notation used in  [GGW22, Thm. 2.1] and Proposition 1.2.
Recall that for all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}$. For all $1\leq k\leq n$, let $\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be the orthogonal projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$. That is:
@@
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$.
@@
Let $1\leq k\leq n$ and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n}}(0,1)$ satisfying
$\mu(B_{\mathfrak{r}}(w,\delta)\cap\Theta)\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
@@
We deduce this from [GGW22, Thm. 2.1]. The argument is more or less standard. Indeed it is similar to the deduction of Theorem 1.1 from Lemma 2.1 and to (a finitary version of) the argument in [GGW22, §2].
Since $1\leq k\leq n$ is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$. Adapting the notation $m^{\delta}$ from the previous section: for every $r\in[\frac{1}{2},1]$ and all $w\in F$, put
@@
First note that, for every $r\in I_{\rm bad}$, the number of $\delta$-boxes $\{\mathsf{B}_{i,r}\}$ required to cover $\mathfrak{p}_{r}(F_{\rm bad}(r))$ is $\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}$. Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n}}(0,1)$; note that
@@
Select a maximal $\delta$-separated subset $\Lambda_{\delta}\subset I_{\rm bad}$ and extend this to a maximal $\delta$-separated subset $\hat{\Lambda}_{\delta}$ of $[\frac{1}{2},1]$. The definition of $F_{\rm bad}$
@@
$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n})\delta^{-1-\alpha+D\eta}$
@@
where $D=10^{10n}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
@@
We now turn to the proof of Proposition 1.2.
Proof of Proposition 1.2.
Let $\delta$ and $k$ be as in Proposition 1.2. As before, we let $\mu$ be the uniform measure on a finite set $F$.
Apply Theorem 3.1 with $\delta^{\prime}:=e^{-t}\delta\geq\delta_{0}$ and with $k$ and $k-1$. Then for every $r\in J_{\delta}$ and $w\in F_{\delta^{\prime},r}$, we have
(3.3) $\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(\ell)}(w)-\mathfrak{p}_{r}^{(\ell)}(w^{\prime})\|\leq e^{-t}\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}(e^{-t}\delta)^{\min(\alpha,\ell)-\varepsilon}$
for $\ell=k,k-1$.
Also note that if $10<\alpha\leq k-$ 1, then
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k-1)}(w)-\mathfrak{p}_{r}^{(k-1)}(w^{\prime})\|\leq e^{-t}\delta\}.$
Thus (3.3), with $\ell=k-1$ implies
(3.4) $\mu(\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\})\ll_{\varepsilon}C_{0}e^{-\alpha t}\delta^{\alpha-\varepsilon}$
as we claimed.
Assume now that $k-1<\alpha\leq k$. Let $Q\subset{\rm Span}\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}\simeq\mathbb{R}^{k}$ denote the box $\delta^{\prime}\times\cdots\times\delta^{\prime}\times\delta$ centered at the origin, where directions correspond directions of $\xi^{(i)}(r)$. That is: $Q$ is box of size $\delta^{\prime}$ in the direction of $\xi^{(i)}(r)$ for all $11\leq i\leq k-$ 1 and of size $\delta$ in the direction of $\xi^{(k)}(r)$. We cover
$Q^{\prime}:=(Q+\pi_{t,r}(w))\cap\pi_{t,r}(F_{\delta^{\prime},r})$
with $\ll e^{t}$ many boxes $B_{i}$ of size $\delta^{\prime}$. Then (3.3), applied with $2\delta^{\prime}$, implies
(3.5) $\mu(Q^{\prime})$ $\ll_{\varepsilon}(C_{0}(2\delta^{\prime})^{\alpha-\varepsilon})\cdot e^{t}$
$\ll_{\varepsilon}C_{0}e^{(1-\alpha)t}\delta^{\alpha-\varepsilon}.$
Finally, note that
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset Q+\pi_{t,r}(w).$
This and (3.5) complete the proof in this case as well. ∎

### v2

@@ 1. Introduction
Indeed we first study the problem in dimension $3$, where an elementary argument inspired by [OV18] will be used. Then we use the main results in [GGW22] to study a similar problem in all dimensions.
Let us fix some notation in order to state the results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
@@
The following is one of the results of this paper.
@@
Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set and let $\mu$ denote the uniform measure on $F$. Assume that
$\mu(B_{\mathbb{R}^{3}}(w,\delta))\leq C_{0}\cdot\delta^{\alpha}\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
@@
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset there exists $\mathcal{E}=\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|\mathcal{E}|\ll_{\varepsilon}\delta^{\star\varepsilon}$ so that the following holds.
Let $r\in[0,1]\setminus\mathcal{E}$, then there exists $F_{t,\delta,r}\subset F$ with
$\mu(F\setminus F_{t,\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}$
such that for all $w\in F_{t,\delta,r}$, we have
$\mu\Bigl{(}\{w^{\prime}\in F_{t,\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}$
@@
As mentioned earlier, the proof of Theorem 1.1 is based on elementary arguments inspired by [OV18].
However, this approach does not readily extend to higher dimensions, nor does it provide a clear path to overcoming the barrier $\alpha\leq 3/2$ in the context of Theorem 1.1. We now present a similar general result on dimension improvement in irreducible ${\rm{SL}}_{2}(\mathbb{R})$-representations. In dimensions $3,5,7$, and $11$ results of this nature drive the Margulis function estimate in [LMWY25], see also [LMW22]. The proof relies on deep ingredients. In particular, our argument makes crucial use of [GGW22, Thm. 2.1].
Let $V$ be an irreducible representation of ${\rm{SL}}_{2}(\mathbb{R})$ with $\dim V=n+1$. We fix a norm on $V$. Note that since all norm on a finite dimensional vector space are equivalent, different choice of norms will only create a constant factor in the dimension estimate. Let
$u_{r}=\begin{pmatrix}1&r\\ &1\end{pmatrix}$
and
$a_{t}=\begin{pmatrix}e^{{t}/2}&\\ &e^{-{t}/2}\end{pmatrix}.$
We have the following result on dimension improvement under the push by $a_{t}u_{r}$ for a discretized measure in an irreducible representation of ${\rm{SL}}_{2}(\mathbb{R})$.
1.2 Theorem.
Let $\alpha\in(0,n+1)$, $C\geq 1$ and $0<\delta_{0}\leq 1$. Let $F$ be a finite set in $B_{V}(0,1)$ and let $\mu$ denote the uniform measure on $F$. Assume that
$\mu(B_{V}(x,\delta))\leq C\delta^{\alpha}$
for all $\delta\in[\delta_{0},1]$. Then the following holds for all $\epsilon\in(0,\frac{1}{10^{4}n}\min\{\alpha,n+1-\alpha\})$.
For all $\delta\in[\delta_{0},\frac{1}{100}]$ and ${t}\gg_{\epsilon}1$ with $e^{\frac{n}{2}{t}}\delta_{0}\leq\delta\leq e^{-\frac{n}{2}{t}}$, there exists $\mathcal{E}=\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|\mathcal{E}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ satisfying the following.
For all $r\in[0,1]\setminus\mathcal{E}$, there exists $F_{t,\delta,r}\subset F$ with $\mu(F\setminus F_{t,\delta,r})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in F_{t,\delta,r}$, we have
$\mu(\{w^{\prime}\in F_{t,\delta,r}:\|a_{t}u_{r}(w^{\prime})-a_{t}u_{r}(w)\|\leq\delta\})\ll_{\epsilon}Ce^{-\frac{\varpi(\alpha){t}}{2}}\delta^{\alpha-\epsilon}$
where
$\varpi(\alpha)=\max\{n\alpha-\lfloor\alpha\rfloor(\lfloor\alpha\rfloor+1),\lfloor\alpha\rfloor(2n-\lfloor\alpha\rfloor+1)-n\alpha\}.$
All the implicit constant here depends only on $n$.
Remark.
Note that if $\alpha$ varies in compact subset of $(0,n+1)$, $\varpi(\alpha)$ is uniformly bounded away from 00.
Remark.
When $n=2$, the representation is the adjoint representation of ${\rm{SL}}_{2}(\mathbb{R})$. We have
$\frac{1}{2}\varpi(\alpha)=\begin{cases}\alpha&0<\alpha\leq 1,\\ \max\{2-\alpha,\alpha-1\}&1<\alpha\leq 2,\\ 3-\alpha&2<\alpha\leq 3.\end{cases}$
See Figure 1.
33331111
Figure 1. Graph of the piecewise function $\varpi(\cdot)$
Remark.
We remark that Theorem 1.2 is a generalization of Theorem 1.1 when ${t}=|\log\delta|$. In this case, the pre-image of a $\delta$-disk under the expanded projections in Theorem 1.1 is of size $\delta^{2}\times\delta\times 1$, which coincide with the pre-image of a $\delta$-ball under $a_{t}u_{r}$.
@@
Recall that $\mu$ denotes the uniform measure on $F$. That is:
@@
3. Proof of Theorem 1.2
We will prove Theorem 1.2 in this section. Before starting the proof let us fix some notation.
For all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!},\tfrac{r^{n+1}}{(n+1)!})\in\mathbb{R}^{n+1}$. For all $11\leq k\leq n+$ 1, let
$\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n+1}\to\mathbb{R}^{k}$
be the projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$ defined by
@@
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n+1}$. We define $\mathfrak{p}^{(0)}_{r}$ to be the zero map for all $r\in[0,1]$.
This family of projections has the following two properties.
  1. (1)
    The projection $\mathfrak{p}^{(n+1)}_{r}$ is exactly the action of $u_{r}$ in this representation.
  2. (2)
    The difference between $\mathfrak{p}_{r}^{(k)}$ and the orthogonal projection to the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$ is a bi-Lipschitz map where the Lipschitz constant depends only on $n$.
Recall that we let
$u_{r}=\begin{pmatrix}1&r\\ &1\end{pmatrix}$
and
$a_{t}=\begin{pmatrix}e^{{t}/2}&\\ &e^{-{t}/2}\end{pmatrix}.$
With this normalization, there exists a basis for the irreducible representation $V$ so that we can identify $V\cong\mathbb{R}^{n+1}$ and for all $w\in\mathbb{R}^{n+1}$,
$u_{r}.w={}$ $\biggl{(}w\cdot\xi^{(1)}(r),\cdots,w\cdot\xi^{(n)}(r)\biggr{)},$
$a_{{t}}.w={}$ $(e^{\frac{n}{2}{t}}w_{1},e^{\frac{n-2}{2}{t}}w_{2}\cdots,e^{-\frac{n}{2}{t}}w_{n+1}).$
For a probability measure $\mu$ supported on $B_{V}(0,1)$, we define the following notions. If $\mu(A)>0$, we let
$\mu_{A}=\tfrac{1}{\mu(A)}\mu|_{A}$
be the normalized restriction of $\mu$ on $A$.
In this section, we will call cubes of the form $\prod_{i=1}^{n+1}[\frac{n_{i}}{2^{k}},\frac{n_{i}+1}{2^{k}})$ dyadic cubes. For a dyadic cube $Q\subset\mathbb{R}^{n+1}$, we set ${\rm{Hom}}_{Q}$ to be the homothety that maps $Q$ to $[0,1)^{n+1}$. We set
$\mu^{Q}=({\rm{Hom}}_{Q})_{\ast}\mu_{Q}$
to be the rescaled normalized restriction of $\mu$ on $Q$. We also refer to $\mu^{Q}$ as the conditional measure of $\mu$ on $Q$. We use the notion $\mathcal{D}_{\rho}$ for the collection of dyadic $\rho$-cubes and $A_{Q}=A\cap Q$.
As it was already mentioned, we will prove Theorem 1.2 using [GGW22, Thm. 2.1]. We record the following consequence of  [GGW22, Thm. 2.1].
@@
Let $11\leq k\leq n+$ 1 and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n+1}}(0,1)$ satisfying
$\mu(B_{\mathbb{R}^{n+1}}(w,\delta))\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
@@
We deduce this from [GGW22, Thm. 2.1]. The argument is more or less standard. Indeed it is similar to the deduction of Theorem 1.1 from Lemma 2.1 and to (a finitary version of) the argument in [GGW22, §2]. See also [JL24, Section 2].
Since $11\leq k\leq n+$ 1 is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$. Adapting the notation $m^{\delta}$ from the previous section: for every $r\in[\frac{1}{2},1]$ and all $w\in F$, put
@@
First note that, for every $r\in I_{\rm bad}$, the number of $\delta$-boxes $\{\mathsf{B}_{i,r}\}$ required to cover $\mathfrak{p}_{r}(F_{\rm bad}(r))$ is $\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}$. Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n+1}}(0,1)$; note that
@@
Select a maximal $\delta$-separated subset $\Lambda_{\delta}\subset I_{\rm bad}$ and extend this to a maximal $\delta$-separated subset $\hat{\Lambda}_{\delta}$ of $[\frac{1}{2},1]$.
@@
$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n+1})\delta^{-1-\alpha+D\eta}$
@@
where $D=10^{10(n+1)}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n+1})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
@@
Now we prove Theorem 1.2.
Proof.
Without loss of generality, we assume both $e^{\frac{n{t}}{2}}\delta$ and $\delta$ are dyadic numbers. This will only result a constant factor depending only on the ambient space in the final estimate.
For simplicity, let $\rho=e^{\frac{n{t}}{2}}\delta$ and $k=\lfloor\alpha\rfloor$. Fix $\delta$ satisfying $e^{\frac{n}{2}{t}}\delta_{0}\leq\delta\leq e^{-\frac{n}{2}{t}}$.
Let $P$ be the tube of size
$e^{-n{t}}\times\cdots\times e^{-{t}}\times 1$
Centered at the origin, with its sides aligned along the coordinate axes, via the identification $V\cong\mathbb{R}^{n+1}$. We note that $u_{r}^{-1}a_{{t}}^{-1}B(x,\delta)$ is a translate of $u_{r}^{-1}.P$. We denote it by $w+u_{r}^{-1}.P$.
We will prove the theorem by estimating $\mu(w+u_{r}^{-1}.P)$. First note that since tubes $w+u_{r}^{-1}.P$ lies in $O(1)$ many dyadic $\rho$-cubes, it suffices to estimate $\mu_{Q}(w+u_{r}^{-1}.P)$ for those dyadic $\rho$-cubes $Q$ intersecting $w+u_{r}^{-1}.P$ non-trivially. To this end, we first study the dimension condition for those conditional measure $\mu^{Q}$.
For all $s\geq\rho^{-1}\delta_{0}$, we have
$\mu^{Q}(B(x,s))={}$ $\mu_{Q}({\rm{Hom}}_{Q}^{-1}B(x,s))$
[…]
Now we study the relation between $\mu_{Q}(w+u_{r}^{-1}.P)$ with the projections $\Bigl{\{}\mathfrak{p}_{r}^{(l)}\Bigr{\}}_{r,l}$. This follows from the standard ‘projection and slicing’ picture, as we now explicate . Namely, in our case, for all $l=0,\cdots,n+1$, there exists some rectangle $R_{l}(w)\subset\mathbb{R}^{l}$ with sides parallel to the axis and side-length
$e^{-n{t}}\times\cdots\times e^{-(n-l+1){t}}$
so that
$w+u_{r}^{-1}.P\subseteq\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}^{-1}(R_{l}(w)).$
It suffices to study $\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}_{\ast}\mu^{Q}(R_{l}(w))$. We will focus on the cases where $l=k$ and $l=k+1$.
Apply Theorem 3.1 with the measures $\{\mu^{Q}\}_{Q}$, $\mathfrak{p}^{(k)}_{r}$, scale $e^{-n{t}}$, and $\epsilon/n$. Using Fubini’s theorem, there exists $\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|[0,1]\setminus\mathcal{E}_{{t},\delta}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $r\in[0,1]\setminus\mathcal{E}_{{t},\delta}$, there exists $\mathcal{D}_{\rho}(r)$ with
$\sum_{Q\notin\mathcal{D}_{\rho}(r)}\mu(Q)\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
Moreover, for all $Q\in\mathcal{D}_{\rho}(r)$, there exists $F_{Q}(r)\subset F_{Q}$ with $\mu_{Q}(F_{Q}\setminus F_{Q}(r))\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in V$, we have
$\mu^{Q}\Bigl{(}F_{Q}(r)\cap(w+u_{r}^{-1}.P)\Bigr{)}\leq{}$ $\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}_{\ast}\mu^{Q}(R_{k}(w))$
$\ll_{\epsilon}{}$ $\frac{C\rho^{\alpha}}{\mu(Q)}\prod_{j=0}^{k-1}e^{-(n-j){t}}e^{\epsilon{t}}.$
The last inequality follows from the fact that we can cover $R_{k}(w)$ with $O(\prod_{j=0}^{k-1}e^{j{t}})$ many $e^{-n{t}}$ balls.
Let $F_{r}^{(k)}=\sqcup_{Q\in\mathcal{D}_{\rho}(r)}F_{Q}(r)$, we have
[…]
Now we apply Theorem 3.1 with $\{\mu^{Q}\}_{Q}$, $\mathfrak{p}^{(k+1)}_{r}$, scale $e^{-n{t}}$, and $\epsilon/n$. Using Fubini’s theorem, again, there exists $\mathcal{E}_{{t},\delta}^{\prime}\subset[0,1]$ with $|[0,1]\setminus\mathcal{E}_{{t},\delta}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $r\in[0,1]\setminus\mathcal{E}_{{t},\delta}^{\prime}$, there exists $\mathcal{D}_{\rho}(r)^{\prime}$ with
$\sum_{Q\notin\mathcal{D}_{\rho}(r)^{\prime}}\mu(Q)\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
Moreover, for all $Q\in\mathcal{D}_{\rho}(r)^{\prime}$, there exists $F_{Q}(r)^{\prime}\subset F_{Q}$ with $\mu_{Q}(F_{Q}\setminus F_{Q}(r)^{\prime})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in V$, we have
$\mu^{Q}(F_{Q}(r)^{\prime}\cap(w+u_{r}^{-1}.P))\leq{}$ $(\mathfrak{p}_{r}^{(k+1)})_{\ast}\mu^{Q}(R_{k+1}(w))$
$\ll_{\epsilon}{}$ $\frac{C\rho^{\alpha}}{\mu(Q)}\Biggl{(}\prod_{j=0}^{k}e^{j{t}}\Biggr{)}e^{-n\alpha{t}}e^{\epsilon{t}}.$
As before, the last inequality follows from the fact that we can cover $R_{k+1}(w)$ via $O(\prod_{j=0}^{k}e^{j{t}})$ many $e^{-n{t}}$ balls.
Let $F_{r}^{(k+1)}=\sqcup_{Q\in\mathcal{D}_{\rho}(r)}F_{Q}(r)$, we have
[…]
Let $\mathcal{E}=\mathcal{E}_{{t},\delta}\cup\mathcal{E}_{{t},\delta}^{\prime}$ and for all $r\notin\mathcal{E}$, let $F_{t,\delta,r}=F_{r}^{(k)}\cap F_{r}^{(k+1)}$. We have
$|[0,1]\setminus\mathcal{E}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$
and
$\mu(F\setminus F_{r})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
Combining (3.3) and (3.4), the theorem follows. ∎